# Axion Astrophysics

### v1

@@ 5.2 Horizontal branch stars: the $R$, $R_{1}$ and $R_{2}$ parameters
As a further test, Ref. [33] repeated the analysis under to opposite and extreme assumptions about the original He content, namely: $Y=0.2472\pm 0.0003$, corresponding to the primordial He mass fraction from extant Big-Bang-Nucleosynthesis calculations with effective number of neutrinos $N_{\rm eff}=3.046$ and baryon density from the Planck legacy [334], and $Y=0.269$, that is the solar He mass fraction, as obtained with extant standard solar models [335]. Note that the latter is a very conservative assumption, considering that the Sun is a pop I star, belonging to the galactic disk, and formed about 8 Gyr later than the bulk of the GCs. As a result, it was found $g_{a\gamma}<0.5\times 10^{-10}$ GeV-1 and $g_{a\gamma}<0.76\times 10^{-10}$ GeV-1, for the primordial and the solar He assumptions, respectively.
@@
Figure 35: SCMDs of GCs harboring two stellar populations, having the same age and metallicity, but different He mass fractions: black dots refer to stars with primordial He abundance ($Y=0.248$), while red dots represent stars with enhanced He. The corresponding values of $R$ are reported in each panel. Note that $R=1.408$ is found in case of a single stellar population with primordial He.
@@ 6.1 Structure and evolution of white dwarfs
At very early times, the main contributor to the luminosity of the nascent WD is hydrogen burning via CNO cycle and is during this epoch that the final thickness of the hydrogen layer is configured. This stage is very short, $\sim 10^{4}$ yr, nuclear reactions cease abruptly and the luminosity drops from $\log(L/L_{\odot})\approx 4$ to $\log(L/L_{\odot})\approx 1-2$ [355]. From this moment neutrino emission becomes dominant, Fig. 38, and forces the different thermal structures to converge to a unique one for models dimmer than $\log(L/L_{\odot})\approx-1.5$ [361]. The comparison between the evolution times obtained with the LPCODE and the BaSTI codes using exactly the same inputs differ at this epoch by an 8% as a consequence of the different converged models at the beginning of the cooling sequence [362]. This means that hot WDs have to be used with care when testing new physics.
After $10^{7}-10^{8}$ yr, depending on the mass of the star, neutrino emission rapidly drops at this stage because of its steep dependence on the core temperature, $L_{\nu}\propto T^{7}_{C}$, and photon losses become dominant. At this epoch photon luminosity is controlled by a thick non-degenerate layer with an opacity dominated by hydrogen, if present, and helium that is weakly dependent of the metal content [363, 361, 355]. At the same time, the core of the WD behaves like a fluid that can be described as a Coulomb plasma not very strongly, $\Gamma<178$, coupled [364, 365, 366] [^17] The Coulomb parameter $\Gamma=Z^{2}e^{2}/akT$, where $a=[3/(4\pi n_{\rm i})]^{1/3}$ is the ion-sphere radius, $n_{\rm i}$ is the ion number density and $k$ is the Boltzmann constant, compares the energy of the Coulomb interaction with the thermal energy. In the case of one component plasma the transition liquid-solid occurs for $\Gamma\simeq 178$. A key ingredient is the electron conductivity at the frontier between moderate and strong degeneracy. Models predict longer cooling times at high luminosities and shorter cooling times at low luminosities [367] when the Blouin et al. conductivities [368] are used instead of those of Cassisi et al[369].
@@
Besides the phenomenological differences, the must important property is that DAs cool down more slowly than non-DAs since hydrogen is more opaque than helium, Fig. 41. Therefore, the nature of the envelope plays a very important role during the cooling stage dominated by photons. At present there are several properties of WDs that allow them to be used as physical laboratories: i) the secular drift of the period of pulsation of variable WDs, and ii) the shape of the luminosity function.
@@ 6.2 Axion bounds from the secular drift of the pulsation period
Figure 44: Values of the axion-electron coupling constant $g_{ae}$ obtained form the analysis of variable WDs.
@@
As shown in Fig. 43, measuring the secular drift is difficult and the observed values have been evolving with time, being $(5.12\pm 0.82)\times 10^{-15}$ ss-1 the most recent value [410]. At the same time, modeling variable WDs has become more and more robust [416, 417, 418, 415] and the predicted values have also evolved. At present the predicted one is $(1.25\pm 0.09)\times 10^{-15}$ ss-1 [410]. Consequently, the axion-electron coupling constant necessary to account for observations has also evolved. Fig. 44 shows this evolution in the G117-B15A case, the last one being $g_{ae}=(5.66\pm 0.57)\times 10^{-13}$ ($m_{a}\cos^{2}\beta=20\pm 2$ meV for DFSZ axions) [410].
The secular drifts of R548 [411] and L19-2 [412] are displayed in Fig. 43. In the second case, the drift was measured for two periods, 113 and 192 seconds, and the same value of $\dot{P}=(3.0\pm 0.6)\times 10^{-15}$ ss-1 was obtained. The values of $g_{ae}$ necessary to account for the discrepancies between the observed and predicted drifts are displayed in Fig. 44 [417, 419, 420, 421].
A similar analysis has been performed with PG 1351 + 489, a DB variable with a drift of $\dot{P}=(2.0\pm 0.69)\times 10^{-13}$ ss-1 for a period of 489 s, the one with the largest amplitude [422]. The theoretical drift is expected to be $\dot{P}=(0.81\pm 0.5)\times 10^{-15}$ ss-1 [421], for which reason the axion-electron coupling constant necessary to fit the observed value should be $g_{ae}<5.5\times 10^{-13}$ [423], a value that is concordant with those obtained with the DAVs (Fig. 44). Notice that this WD still has a strong neutrino emission and the result could be influenced by an extra emission if neutrinos have a magnetic moment.
@@ 6.3 Axion bounds from the luminosity function
since written in this way it avoids not only he numerical derivative of the IFMR but also that of the cooling time in order to obtain the characteristic cooling time [433]. In deriving this equation it has been assumed that WDs are not destroyed nor created via merging of binaries and that the ensemble is closed, i.e. there is not an important flux of stars in and out of the ensemble under consideration.
@@
The first WDLF was derived by Weidemann [434]. Fig. 45 (panel A) shows the state of the art by the end of the 90s, before the advent of the large cosmological surveys. The monotonic growing of the number of stars with the magnitude strongly supported the idea that the evolution of WDs was just a cooling process and, consequently, that the sudden cut-off was caused by the finite age of the Galaxy [435] but the large dispersion of bins prevented the determination of the slope, and the unique tools to constrain physics were the expected populations of the bins and the position of the cut-off [436]. With the advent of the large cosmological surveys, like the SDSS and the SCSS catalogues, the size of the samples increased to several thousand of stars thus allowing the use of the slope as a tool for testing new physics, Fig. 45, panel B. For instance, the change of the slope in this figure shows the change from neutrino to photon dominated cooling. At present, thanks to the accuracy of the photometric and astrometric data provided by the Gaia mission, it has been possible obtain the WDLFs for a statistically complete 100 pc sample, Fig. 45 (panel C).
@@
Bright WDs have had no time to cool down, $t_{\rm cool}$ is small, and, as a consequence of the strong dependence of the MS lifetimes with mass, $M_{i}$ is almost constant and independent of the luminosities under consideration. As a consequence, the slope essentially depends on the averaged characteristic cooling time and only weakly on the star formation history. This average is dominated by low-mass WDs and, as far as the mass distribution is not strongly perturbed by the adopted SFR or initial mass function, is representative of the intrinsic properties of WDs. Notice that this is not true if recent star formation bursts are present since, in this case, low-mass MS stars have no time to become WDs and $M_{i}$ becomes dependent on the luminosity under consideration. On the contrary, if the burst is old enough, the slope of the bright region is maintained and differences appear at larger magnitudes. See Fig. 46, solid lines. Another important property is that in this bright region the slope of the relationship between the luminosity and the core temperature of DA and non-DA WDs almost coincide and both luminosity functions almost overlap in this luminosity interval after normalization [437].
@@
In the case of axions, this technique was used for the first time using a preliminary form of Fig. 45-B [430] obtaining $g_{ae}=(1.4^{+0.9}_{-0.8})\times 10^{-13}$ [438]. This result was reexamined by Ref. [439] using all the luminosity functions of Fig. 45-B but using in this case a consolidated set of data, and a self consistent treatment of the neutrino cooling, concluding that the existence of DFSZ axions with a coupling constant $g_{ae}$ in the range of $(0.7-2.1)\times 10^{-13}$ could improve the agreement between theoretical and observed WDLFs, Fig. 47.
Figure 47: Axion-electron couplings obtained from different luminosity functions. See text for details.
Given the degeneracy between the stellar and galactic terms of Eq. (123) it is natural to wonder if the shape of the WDLF can be attributed to axions or if it is just a consequence of changes in the SFR like those introduced by recent bursts of star formation, solid red line in Fig. 46.
@@
    The discrepancies between the theoretical calculations and the observed luminosity functions of the thin and thick disks and halo [429] decrease if axions with $g_{ae}$ in the range of $(2.24-4.48)\times 10^{-13}$ are introduced [440] [^22] There is some tension between the halo and disk results, but the $2\sigma$ bounds are compatible., Fig. 47, label RH.
@@
    Munn et al[441] improved the disk luminosity function assuming a constant scale height above the Galactic plane and recomputed that of the halo. As before, the inclusion of axions improves the concordance between theory and observations both in the disk and in the halo. The best it was obtained for $g_{ae}\approx 2.24\times 10^{-13}$, Fig. 47, label M.
@@
    Kilic et al[442] reexamined the Munn loc.cit. data assuming different variable scale heights above the Galactic plane. Although this improved the results, the concordance was better when axions with a $g_{ae}\approx 2.24\times 10^{-13}$ were included, Fig. 47, label K.
@@
Another possibility is provided by massive WDs. Since the lifetime of their progenitors is short compared with the cooling time, their luminosity function closely follows the temporal variation of the SFR, dashed lines of Fig. 46 and in this case it is possible to directly obtain the star formation rate $\Psi$ [443]. Recently, the WDLF for stars with a mass in the range of 0.9-1.1 M and within a distance of 100 pc has been obtained [444]. This luminosity function presents a bump that has been associated to the process of crystallization, however, because of the degeneracy of Eq. (123), it is always possible to find a SFR able to reproduce the observed WDLF.
@@
Figure 48: SFRs per unit volume obtained assuming solidification and axion cooling with $g_{ae}=0.0,1.12,2.24,and4.48~{}\times 10^{-13}$ (black, red, blue and magenta solid lines). Dotted dashed lines, SFRs obtained from MS stars [445] divided by an arbitrary scale height constant and age dependent respectively (see text for details).
Figure 49: Luminosity functions resulting from the SFRs displayed in Fig. 48 normalized to $M_{\rm bol}=12$ of that of WDs within a distance of 100 pc.
@@
The black solid line represents the SFR obtained from the luminosity function of massive white dwarfs using the old BaSTI models [362] which include crystallization and the associated separation of carbon and oxygen for white dwarfs with solar metallicity. As it can be seen the SFR increases with the age and reaches a maximum around 7-8 Gyr ago and decreases at larger ages. A noticeable feature is the existence of a prominent peak around 2.5 Gyr ago, being the exact position and height model depending (see Refs. [443, 446] for details). It is important to notice the the SFR obtained in this way does not takes into account the secular evolution of the sample caused by radial migrations, scale height inflation or merging of double degenerate stars. The peak appearing around 0.2 Gyr is in the limit of applicability of the method and deserves more attention.
@@
The Gaia data release 2 has allowed to obtain the SFR per unit of disk surface for MS stars with $G<12$ and the Besançon Galaxy Model [445]. This rate was divided by an arbitrary and constant scale height above the galactic plane in order to compare the SFR per unit of surface of the Galactic with the SFR per unit volume obtained from white dwarfs. Both methods predict a concordant burst of star formation $\sim 2.5$ Gyrs ago but diverge at large ages, Fig. 48. This tension may have several origins: a local delay in starting the star formation process, a different behavior of the outer and inner regions of the disk, a vertical dilution larger than expected caused by an early collision or just an efficient conversion of DA WDs into non-DAs [443].
@@
If axions are included, red, blue and magenta solid lines in Fig. 48, the bump moves towards shorter ages and values $g_{ae}\gtrsim 4\times 10^{-13}$ are not compatible with the results obtained with Gaia. However, if a time dependent height scale, $h_{\rm old}>h_{\rm young}$, dashed line, is adopted [447], the burst moves to shorter ages and values of $g_{ae}\sim 2.24\times 10^{-13}$ are acceptable while values $g_{ae}\gtrsim 4.5\times 10^{-13}$ can be rejected, Fig. 47  [^23] Notice that these results are very sensitive to the evolution of the DA/non-DA ratio and to the contribution of binary mergers. [446]. Fig. 49 displays the luminosity functions built with the SFRs obtained from massive WDs assuming different axion masses represented in Fig. 48. The deficit of WDs predicted by theoretical models around $M_{\rm bol}=10$ may have different origins, the dependence on the initial models, which are poorly known, the contribution of mergers, or a very recent burst of star formation.

### v2

@@ 5.2 Horizontal branch stars: the $R$, $R_{1}$ and $R_{2}$ parameters
As a further test, Ref. [33] repeated the analysis under two opposite and extreme assumptions about the original He content, namely: $Y=0.2472\pm 0.0003$, corresponding to the primordial He mass fraction from extant Big-Bang-Nucleosynthesis calculations with effective number of neutrinos $N_{\rm eff}=3.046$ and baryon density from the Planck legacy [341], and $Y=0.269$, that is the solar He mass fraction, as obtained with extant standard solar models [342]. Note that the latter is a very conservative assumption, considering that the Sun is a pop I star, belonging to the galactic disk, and formed about 8 Gyr later than the bulk of the GCs. As a result, it was found $g_{a\gamma}<0.5\times 10^{-10}$ GeV-1 and $g_{a\gamma}<0.76\times 10^{-10}$ GeV-1, for the primordial and the solar He assumptions, respectively.
@@
Figure 35: SCMDs of GCs harboring two stellar populations, having the same age and metallicity, but different He mass fractions: black dots refer to stars with primordial He abundance ($Y=0.248$), while red dots represent stars with enhanced He. The corresponding values of $R$ are reported in each panel. Note that $R=1.408$ is found in the case of a single stellar population with primordial He.
@@ 6.1 Structure and evolution of white dwarfs
At very early times, the main contributor to the luminosity of the nascent WD is hydrogen burning via CNO cycle and it is during this epoch that the final thickness of the hydrogen layer is configured. This stage is very short, $\sim 10^{4}$ yr, nuclear reactions cease abruptly and the luminosity drops from $\log(L/L_{\odot})\approx 4$ to $\log(L/L_{\odot})\approx 1-2$ [362]. From this moment neutrino emission becomes dominant, Fig. 38, and forces the different thermal structures to converge to a unique one for models dimmer than $\log(L/L_{\odot})\approx-1.5$ [368]. The comparison between the evolution times obtained with the LPCODE and the BaSTI codes using exactly the same inputs differ at this epoch by an 8% as a consequence of the different converged models at the beginning of the cooling sequence [369]. This means that hot WDs have to be used with care when testing new physics.
After $10^{7}-10^{8}$ yr, depending on the mass of the star, neutrino emission rapidly drops at this stage because of its steep dependence on the core temperature, $L_{\nu}\propto T^{7}_{C}$, and photon losses become dominant. At this epoch photon luminosity is controlled by a thick non-degenerate layer with an opacity dominated by hydrogen, if present, and helium that is weakly dependent of the metal content [370, 368, 362]. At the same time, the core of the WD behaves like a fluid that can be described as a Coulomb plasma not very strongly coupled, $\Gamma<178$ [371, 372, 373] [^20] The Coulomb parameter $\Gamma=Z^{2}e^{2}/akT$, where $a=[3/(4\pi n_{\rm i})]^{1/3}$ is the ion-sphere radius, $n_{\rm i}$ is the ion number density and $k$ is the Boltzmann constant, compares the energy of the Coulomb interaction with the thermal energy. In the case of one component plasma the transition liquid-solid occurs for $\Gamma\simeq 178$. A key ingredient is the electron conductivity at the frontier between moderate and strong degeneracy. Models predict longer cooling times at high luminosities and shorter cooling times at low luminosities [374] when the Blouin et al. conductivities [375] are used instead of those of Cassisi et al[376].
@@
Besides the phenomenological differences, the most important property is that DAs cool down more slowly than non-DAs since hydrogen is more opaque than helium, Fig. 41. Therefore, the nature of the envelope plays a very important role during the cooling stage dominated by photons. At present there are several properties of WDs that allow them to be used as physical laboratories: i) the secular drift of the period of pulsation of variable WDs, and ii) the shape of the luminosity function.
@@ 6.2 Axion bounds from the secular drift of the pulsation period
Figure 44: Evolution of the axion-electron constant $g_{ae}$ obtained from WD variables. The amplitude of the segments represent the $1\,\sigma$ uncertainty introduced by observations and models. Dots represent the values of $g_{ae}$ that fit the observations if axions alone were responsible of the extra cooling.
@@
Measuring the secular drift is difficult and the observed values have been evolving with time. In the case of G117-B15A the last value that has been obtained is $(5.12\pm 0.82)\times 10^{-15}$ ss-1 [417], as shown in Fig. 43. The codes that have been used to analyse these data are the La Plata Code (LPCODE), which computes model for the ZZ Ceti white dwarf from the appropriate parent star [422, 423], and the White Dwarf Evolution Code (WDEC), which parametrizes the chemical structure of the WD until the optimal model is obtained [424]. As a consequence of the improving of both, theoretical and observational results, the value of $g_{ae}$ necessary to account for the pulsation drift of G117-B15 A via axions has been changing with time (Fig. 44[425, 426, 427, 422]. At present the predicted value is $g_{ae}=(5.66\pm 0.57)\times 10^{-13}$ ($m_{a}\cos^{2}\beta=20\pm 2$ meV for DFSZ axions) [417] (Fig. 44).
The secular drift of R548, also known as ZZ Ceti, is $(3.3\pm 1.1)\times 10^{-15}$ ss-1 [418]. The theoretical analysis performed with WDEC provided an upper limit of $g_{ae}\lesssim 4.8\times 10^{-13}$ ($m_{a}\cos^{2}\beta\lesssim 17$ meV for DFSZ axions), and $g_{ae}=2.8\times 10^{-13}$ ($m_{a}\cos^{2}\beta=10$ meV for the exact fit [428]. The analysis with the LPCODE provided a value $g_{ae}=4.79^{+1.2}_{-1.7}\times 10^{-13}$  [429] [^25] Notice that the values of $g_{ae}$ quoted here were obtained from the observations and the code versions at that moment and that both have been changing with time..
L19-2 has been observed from 1976 to 2012 and presents five independent pulsation modes of which two, the 113 and 192 seconds ones, are non-trapped in contrast with the G117-B15A and R548 cases and have the same pulsation drift, $\dot{P}=(3.0\pm 0.6)\times 10^{-15}$ ss-1 [419] (Fig. 43). The values of $g_{ae}$ necessary to account for the discrepancies between the observed and the predicted drifts of the LPCODE are $g_{ae}=2.8^{+4.2}_{-1.4}\times 10^{-13}$ ($m_{a}=10^{+15}_{-5}$ meV) for the 113 period, and $g_{ae}\lesssim 5\times 10^{-13}$ ($m_{a}\cos^{2}\beta\lesssim 18$ meV) and $g_{ae}\approx 2.52\times 10^{-13}$ ($m_{a}\cos^{2}\beta\approx 9$ meV) for the 192 s one (Fig. 44[430].
The analysis performed with the WDEC code [431] suggests that L19-2 is massive, hot, with a core very rich in oxygen and with a relatively thick H-layer. Concerning pulsation periods, the predicted secular drifts are smaller than the observed ones and not entering in contradiction with the existence of axions
A similar analysis has been performed with PG 1351 + 489, a DB variable with a drift of $\dot{P}=(2.0\pm 0.69)\times 10^{-13}$ ss-1 for a period of 489 s, the one with the largest amplitude [432]. The theoretical drift is expected to be $\dot{P}=(0.81\pm 0.5)\times 10^{-15}$ ss-1 [430], for which reason the axion-electron coupling constant necessary to fit the observed value is $g_{ae}\approx 2.24\times 10^{-13}$ ($m_{a}\cos^{2}\beta\approx 8$ meV), and an upper limit of $g_{ae}<5.5\times 10^{-13}$ ($m_{a}\cos^{2}\beta\approx 8$ meV) [433], a value that is concordant with those obtained with the DAVs (Fig. 44). Notice that this WD still has a strong neutrino emission and the result could be influenced by an extra emission if neutrinos have a magnetic moment.
@@ 6.3 Axion bounds from the luminosity function
since written in this way it avoids not only he numerical derivative of the IFMR but also that of the cooling time necessary to obtain the characteristic cooling time [443]. In deriving this equation it has been assumed that WDs are not destroyed nor created via merging of binaries and that the ensemble is closed, i.e. there is not an important flux of stars in and out of the ensemble under consideration.
@@
The first WDLF was derived by Weidemann [444]. Fig. 45 (panel A) shows the state of the art by the end of the 90s, before the advent of the large cosmological surveys. The monotonic growing of the number of stars with the magnitude strongly supported the idea that the evolution of WDs was just a cooling process and, consequently, that the sudden cut-off was caused by the finite age of the Galaxy [445] but the large dispersion of bins prevented the determination of the slope, and the unique tools to constrain physics were the expected population at each bin and the position of the cut-off [446]. With the advent of the large cosmological surveys, like the SDSS and the SCSS catalogues, the size of the samples increased to several thousand of stars thus allowing the use of the slope as a tool for testing new physics, Fig. 45, panel B. For instance, the change of the slope in this figure shows the change from a cooling dominated by neutrinos to one dominated by photons. At present, thanks to the accuracy of the photometric and astrometric data provided by the Gaia mission, it has been possible to obtain the WDLF for a statistically complete 100 pc sample, Fig. 45 (panel C).
@@
Bright WDs have had no time to cool down, i.e. $t_{\rm cool}$ is small, and, as a consequence of the strong dependence of the MS lifetimes with mass, $M_{i}$ is almost constant and independent of the luminosities under consideration. As a consequence, the slope essentially depends on the averaged characteristic cooling time and only weakly on the star formation history. This average is dominated by low-mass WDs and, as far as the mass distribution is not strongly perturbed by the adopted SFR or initial mass function, is representative of the intrinsic properties of WDs. Notice that this is not true if recent star formation bursts are present since, in this case, low-mass MS stars have no time to become WDs and $M_{i}$ becomes dependent on the luminosity under consideration. On the contrary, if the burst is old enough, the slope of the bright region is maintained and differences appear at larger magnitudes. See Fig. 46, solid lines. Another important property is that in this bright region the slope of the relationship between the luminosity and the core temperature of DA and non-DA WDs almost coincide and both luminosity functions almost overlap in this luminosity interval after normalization [447].
@@
In the case of axions, this technique was used for the first time by Ref. [448] using a preliminary form of the luminosity function obtained by [440] displayed in Fig. 45-B obtaining $g_{ae}=(1.4^{+0.9}_{-0.8})\times 10^{-13}$ (Fig. 47, label Isern et al. 2008). This result was reexamined by Ref. [449] using all the luminosity functions represented in Fig. 45-B and a self-consistent treatment of the neutrino cooling, concluding that the existence of DFSZ axions with a coupling constant $g_{ae}\sim 1.4\times 10^{-13}$ could improve the agreement between theoretical and observed WDLFs and values $g_{ae}\gtrsim 2.8\times 10^{-13}$ could be excluded (Fig. 47, label Miller Bertolami et al. 2014).
Figure 47: Values of the axion-electron coupling constant $g_{ae}$ obtained from different white dwarf luminosity functions. The amplitude of the segments roughly represents the values compatible with the observations taking into account the present uncertainties. Dots represent the values that better fit the observations. See the text for details.
Given the degeneracy between the stellar and galactic terms of Eq. (123) it is natural to wonder if the shape of the WDLF can be attributed to axions or if it is just a consequence of changes in the SFR like those introduced by recent bursts of star formation (solid red line in Fig. 46).
@@
[…]
@@
    Munn et al[451] improved the disk luminosity function assuming a constant scale height above the Galactic plane and recomputed that of the halo. As before, the inclusion of axions improved the concordance between theory and observations both in the disk and in the halo. The best it was obtained for $g_{ae}\approx 2.24\times 10^{-13}$ (Fig. 47, label M) [450].
@@
    Kilic et al[452] reexamined the Munn loc.cit. data assuming different variable scale heights above the Galactic plane. Although this improved the results, the concordance observed and computed luminosity function was better when axions with a $g_{ae}\approx 2.24\times 10^{-13}$ were included (Fig. 47, label K) [450].
@@
Another possibility to break the degeneracy between stellar and galactic properties is provided by massive WDs. Since the lifetime of their progenitors is short compared with the cooling time, their luminosity function closely follows the temporal variation of the SFR, dashed lines of Fig. 46 and in this case it is possible to directly obtain the star formation rate $\Psi$ [453]. The luminosity function for WDs with a mass in the range of 0.9-1.1 M and within a distance of 100 pc has been recently obtained by [454]. This WDLF presents a bump that has been associated to the process of crystallization, but, because of the degeneracy of Eq. (123), it is always possible to find a SFR able to reproduce the observed WDLF.
@@
Figure 48: SFRs per unit volume obtained assuming solidification and axion cooling with $g_{ae}=0.0,1.12,2.24,and4.48~{}\times 10^{-13}$ (black, red, blue and magenta solid lines). Dotted and dashed lines represent the SFRs obtained from MS stars [455] divided, respectively, by a constant and an age dependent scale height (see text for details).
@@
The black solid line of Fig. 48 represents the SFR obtained from the luminosity function of massive white dwarfs using the old BaSTI models [369] which include crystallization and the associated separation of carbon and oxygen for white dwarfs with solar metallicity. As it can be seen the SFR increases with the age and reaches a maximum around 7-8 Gyr ago and decreases at larger ages. A noticeable feature is the existence of a prominent peak around 2.5 Gyr ago, being the its exact position and height model depending (see Refs. [453, 456] for details). It is important to notice the the SFR obtained in this way does not takes into account the secular evolution of the sample caused by radial migrations, scale height inflation or merging of double degenerate stars. The peak appearing around 0.2 Gyr is in the limit of applicability of the method and deserves more attention.
@@
The Gaia data release 2 together the Besançon Galaxy Model has allowed to obtain the SFR per unit of disk surface for MS stars with $G<12$ [455]. In order to compare this rate with the SFR per unit volume obtained from white dwarfs it is necessary to introduce the galactic disk scale height. If a constant scale height is adopted, both methods predict a concordant burst of star formation at $\sim 2.5$ Gyrs ago but diverge at large ages, Fig. 48. This tension may have several origins: a local delay in starting the star formation process, a different behaviour of the outer and inner regions of the disk, a vertical dilution larger than expected caused by an early collision or just an efficient conversion of DA WDs into non-DAs [453].
@@
If axions are included, red, blue and magenta solid lines in Fig. 48, the bump moves towards shorter ages and values $g_{ae}\gtrsim 4\times 10^{-13}$ are not compatible with the results obtained with Gaia. If a time dependent height scale like that of [457], where $h_{\rm old}>h_{\rm young}$ (dashed line), is adopted, values of $g_{ae}\sim 2.24\times 10^{-13}$ are acceptable while values $g_{ae}\gtrsim 4.5\times 10^{-13}$ can be rejected [^26] Notice that these results are very sensitive to the evolution of the DA/non-DA ratio and to the contribution of binary mergers. [456].
Figure 49: Luminosity functions resulting from the SFRs displayed in Fig. 48 normalized to $M_{\rm bol}=12$ as compared whith that of WDs within a distance of 100 pc [442].
Fig. 49 compares the WDLFs built with the SFRs obtained from massive WDs assuming different axion masses represented in Fig. 48 with the luminosity function of the 100 pc sample [442]. The deficit of WDs predicted by theoretical models around $M_{\rm bol}=10$ may have different origins, the dependence on the initial models, which are poorly known, the contribution of mergers, or the aforementioned very recent burst of star formation [458].
The shortage of white dwarfs in the region $8\lesssim M_{\rm bol}\lesssim 13$ of the luminosity functions of disc and halo points out towards an intrinsic origin, being axions a possibility. Nevertheless, other factors not yet well known could be acting over the white dwarf distribution like migration of stars and inflows/outflows of gas in the solar neighborhood, vertical inflation, merging of stars in binary systems, evolution of the DA/non-DA character and so on. The analysis of the globular cluster 47 Tuc [459, 460] seems not to favor the axion option, although the influence of the presence of a black hole in the center, the hydrogen burning in low metallicity WDs, and the presence of several generations of stars may change this perception. If an intrinsic origin is adopted, the problem to be solved is the apparent inconsistency with the results obtained with variable white dwarfs and the bounds set by red giant stars.